A piece of clay sits 0.65 m from the center of a potter’s Wheel. If the potter spins the wheel at a tangential velocity of 4.0 m/s, what is the magnitude of the centripetal acceleration of the piece of clay on the wheel?
1
Expert's answer
2018-05-24T12:52:08-0400
We can find the magnitude of the centripetal acceleration of the piece of clay on the wheel from the definition of the centripetal acceleration:
a_c=v^2/r,
here, a_c is the magnitude of the centripetal acceleration of the piece of clay on the wheel, v is the tangential velocity of the wheel and r is the distance of the piece of clay from the center of a potter’s wheel.
Then, from this formula we get:
a_c=v^2/r=(4.0 m/s)^2/(0.65 m)=24.6 m/s^2 .